OpenStudy (anonymous):

Jessica is 16 years older than James, 15 years ago jessica was 5 times older than james. how old is Jessica.

OpenStudy (experimentx):

x - y =16 (x-15) = 5(y-15) solve this, x for jessica and y for james

OpenStudy (experimentx):

19 and 35 i guess